OpenStudy (anonymous):

The speed of a passenger train is 6 mph faster than the speed of a freight train. The passenger train travels 320 miles in the same time it takes the freight train to travel 290 miles. Find the speed of each train.

OpenStudy (lilg132):

320/(x + 6) = 290/x 320x = 290(x+6) 320x = 290x + 290(6) 30x = 1740 x = 58 and 64 passanger train is 64 mph and freight train 58